OpenStudy (anonymous):

11x^2+122a^2x^2=11a^2 solve x in terms of a

OpenStudy (anonymous):

it should equal -11a^2

jimthompson5910 (jim_thompson5910):

\[\Large 11x^2+122a^2x^2=11a^2\] \[\Large x^2(11+122a^2)=11a^2\] \[\Large x^2=\frac{11a^2}{11+122a^2}\] \[\Large x=\pm\sqrt{\frac{11a^2}{11+122a^2}}\]